Solve the system of equations given below
X+ 3z=12
-x -2y+z=10
3x+5y+2z=-7

Answer by stanbon(75887)

Solve the system of equations given below
X+ 3z=12
-x -2y+z=10
3x+5y+2z=-7
----------------
Substitute for "x" to get 2 equations in 2 variables::
3z-12 - 2y+z = 10
(-9z+36)+5y+2z = -7
-------------------------------
4z -2y = 22
-7z + 5y = -43
---------------------------
2z - y = 11
-7z + 5y = -43
------------------------
Modify for elimination::
10z - 5y = 55
-7z + 5y = -43
------------------------------
Add and solve for "y"::
3z = 12
z = 4
----
Solve for "y using y = 2z-11
y = 8-11 = -3
-----
Solve for "x" using x = -3z+12
x = -12 + 12
x = 0
-----------------
Cheers,
Stan H.
-------------------